# The separating Noether number of small groups

### v2

@@ 5.2. The binary tetrahedral group
having degrees $4,6,12,12$. In fact $h_{12}^{(1)}$ can be omitted, because we have $h_{12}^{(1)}=-4h_{6}^{2}$. Set $v:=[1,\xi^{2}]^{T}$ and $v^{\prime}:=\xi v$. Then $(x_{1}^{2}+x_{2}^{2})(v)=0=(x_{1}^{2}+x_{2}^{2})(v^{\prime})$, hence $f_{2}(v)=0=f_{2}(v^{\prime})$. It follows that $h_{6}(v)=0=h_{6}(v^{\prime})$. Moreover, $h_{8}(v)=16(\xi^{4})^{2}=16=16(\xi^{8})^{2}=h_{8}(v^{\prime})$. Thus no invariants of degree less than $12$ separate $v$ and $v^{\prime}$. However, they have different $G$-orbit, since $h_{12}^{(2)}(v)=4^{3}(\xi^{4})^{3}=-64$, whereas $h_{12}^{(2)}(v^{\prime})=4^{3}(\xi^{8})^{3}=64$. Thus we proved the inequality $\beta_{\mathrm{sep}}(G,V)\geq 12$, implying in turn $\beta_{\mathrm{sep}}^{K}(G)\geq 12$. On the other hand, by [11, Corollary 3.6] we know that $\beta^{K}(G)=12$, implying the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 12$. ∎
5.3. The symmetric group $\mathrm{S}_{4}$ of degree $4$.
Theorem 5.4.
Proof.
We have $\beta^{K}(\mathrm{S}_{4})=9$ by [14, Example 5.3], implying the inequality $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{4})\leq 9$.
In order to show the reverse inequality consider $V=K^{4}$ endowed with the tensor product of the sign representation of $\mathrm{S}_{4}$ and the standard action of $\mathrm{S}_{4}$ on $K^{4}$ via permuting the coordinates. Thus $\pi\in\mathrm{S}_{4}$ maps the variable $x_{j}$ to $\mathrm{sign}(\pi)x_{\pi(j)}$. Denote by $\sigma_{j}$ ($j=1,2,3,4)$ the elementary symmetric polynomial of degree $j$, and set $\Delta:=\prod_{1\leq i<j\leq 4}(x_{i}-x_{j})$. The algebra $K[V]^{\mathrm{S}_{4}}$ is minimally generated by $\sigma_{1}^{2}$, $\sigma_{2}$, $\sigma_{1}\sigma_{3}$, $\sigma_{4}$, $\sigma_{3}^{2}$, $\sigma_{1}\Delta$, $\sigma_{3}\Delta$ (see [14, Example 5.3]).
We shall show that there exists a $v\in V$ with
$(x^{2}+x+1)(x^{2}-x+2)=x^{4}+2x^{2}+x+2\in K[x],$
5.4. The Pauli group
In this section $G$ stands for the Pauli group $G$ of order $16$. Assume the $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ is generated by the matrices
$\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad\begin{bmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\\ \end{bmatrix},\quad\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix}$
in $\mathrm{GL}(K^{2})$.
Theorem 5.5.
Assume that $K$ has an element of multiplicative order $4$. We have the equality $\beta_{\mathrm{sep}}^{K}(G)=7$.
Proof.
Since $G$ is generated by reflections, by the Sheppard-Todd-Chevalley Theorem (see e.g. [3, Theorem 7.2.1] or [16, Section 3.9.4]), $K[x_{1},x_{2}]^{G}$ is a polynomial algebra with two generators. In fact it is generated by $x_{1}^{4}+x_{2}^{4}$ and $(x_{1}x_{2})^{2}$. Moreover, the factor of $K[x_{1},x_{2}]$ modulo the Hilbert ideal (the ideal generated by $x_{1}^{4}+x_{2}^{4}$ and $(x_{1}x_{2})^{2}$) is isomorphic to the regular representation of $G$ (see [8]). The element $f:=(x_{1}^{4}-x_{2}^{4})x_{1}x_{2}$ has the property that $g\cdot f=-f$ where $g$ is any of the above three generators of $G$. One can easily see that $f$ does not belong to the Hilbert ideal. Take an extra indeterminate $t$ and extend the action of $G$ to $K[x_{1},x_{2},t]$ by setting $g\cdot t=-t$ for any of the above three generators of $G$. Then $ft$ is a degree $7$ $G$-invariant separating $v:=[1,1,1]^{T}$ and $v^{\prime}:=[1,1,-1]^{T}$. We claim that $v$ and $v^{\prime}$ can not be separated by an invariant of degree at most $6$. Suppose for contradiction that $h$ is a multihomogeneous $G$-invariant of degree at most $6$ with $h(v)\neq h(v^{\prime})$ (cf. Lemma 3.5). Then $h$ is homogeneous in $t$. Since $v$ and $v^{\prime}$ have the same $(x_{1},x_{2})$ coordinates, any element of $K[x_{1},x_{2}]$ takes the same value on them, so $h$ has positive degree in $t$. Moreover, since $t^{2}$ is $G$-invariant, we may assume that $h=th_{1}$ for some $h_{1}\in K[x_{1},x_{2}]$, where $h_{1}$ is homogeneous of degree at most $5$. By $G$-invariance of $h$ we must have that $g\cdot h_{1}=-h_{1}$ for any of the above three generators $g$ of $G$. However, this $1$-dimensional representation of $G$ occurs only once as a summand in the factor of $K[x_{1},x_{2}]$ modulo the Hilbert ideal, and we found already one occurrence in the degree $6$ homogeneous component. This contradiction proves the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq 7$. The reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 7$ follows from the equality $\beta^{K}(G)=7$ proved in [14, Example 5.4]. ∎
5.5. The Heisenberg group $\mathrm{H}_{27}$
In this section
Write $V$ for the vector space $K^{3}$ endowed with the representation $\psi$.
Proposition 5.6.
Proof.
Denote by $H$ the subgroup of $G$ generated by $a$ and $c$. The algebra $K[V]^{H}$ is generated by the monomials $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$, $x_{1}x_{2}x_{3}$. It follows that each homogeneous $G$-invariant has degree divisible by $3$. The element $b$ fixes $x_{1}x_{2}x_{3}$, and permutes cyclically $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$. So $K[V]^{G}$ is generated by $x_{1}x_{2}x_{3}$ and the generators of $K[x_{1}^{3},x_{2}^{3},x_{3}^{3}]^{\langle b\rangle}$. One easily deduces that the $\langle b\rangle$-invariants in $K[V]^{H}$ (i.e. the $G$-invariants in $K[V]$) of degree at most $6$ are contained in the subalgebra generated by $f_{1}:=x_{1}x_{2}x_{3}$, $f_{2}:=x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$, $f_{3}:=x_{1}^{3}x_{2}^{3}+x_{1}^{3}x_{3}^{3}+x_{2}^{3}x_{3}^{3}$. If $|K|\neq 4$ then there exists a non-zero $\lambda\in K$ with $\lambda^{3}\neq 1$. Now $f_{1}$, $f_{2}$, and $f_{3}$ are all symmetric in the variables $x_{1},x_{2},x_{3}$, therefore they agree on $v:=[1,\lambda,0]^{T}$ and $v^{\prime}:=[\lambda,1,0]^{T}$. On the other hand, $v$ and $v^{\prime}$ have different $G$-orbits. Indeed, the algebra $K[V]^{G}$ contains also $f_{4}:=x_{1}^{3}x_{2}^{6}+x_{1}^{6}x_{3}^{3}+x_{2}^{3}x_{3}^{6}$. We have $f_{4}(v)=\lambda^{6}$, whereas $f_{4}(v^{\prime})=\lambda^{3}$, so $f_{4}(v)\neq f_{4}(v^{\prime})$. Thus $G\cdot v\neq G\cdot v^{\prime}$, and $v$, $v^{\prime}$ can not be separated by invariants of degree $<9$.
Theorem 5.7.
Proof.
By [10, Corollary 18] we have $\beta^{K}(G)=9$, therefore the result follows by Proposition 5.6. ∎
5.6. The non-abelian group $(\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4}$
In this section
$G:=\langle a,b,c\mid a^{2}=b^{2}=c^{4}=1,\quad ab=ba,\quad cac^{-1}=b,\quad cbc^{-1}=a\rangle.$
Theorem 5.8.
Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then we have the equality $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4})=6$.
Proof.
The equality $\beta^{K}(G)=6$ is proved in [15, Proposition 3.10], so it is sufficient to prove that $\beta_{\mathrm{sep}}^{K}(G)\geq 6$. The element $c^{2}$ belongs to the center of $G$, and the factor group $G/\langle c^{2}\rangle$ is isomorphic to the dihedral group $\mathrm{D}_{8}$. The representation
$a\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ -1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix}$
has $c^{2}$ in its kernel and gives the defining representation of $\mathrm{D}_{8}$ as the group of symmetries of a square in the euclidean plane. The element $ab$ also belongs to the center of $G$, and the factor group $G/\langle ab\rangle\cong\mathrm{C}_{4}\times\mathrm{C}_{2}$ is generated by the coset of $c$ and the coset of $a$. Thus $G$ has the $1$-dimensional representations
$a\mapsto 1,\quad b\mapsto 1,\quad c\mapsto\mathrm{i}\qquad\text{ and }\qquad a\mapsto-1,\quad b\mapsto-1,\quad c\mapsto\mathrm{i}.$
Take the direct sum of the above $2$-dimensional and $1$-dimensional representations, and denote by $x_{1},x_{2}$ the standard coordinate functions on the $2$-dimensional summand, and by $t_{1},t_{2}$ the coordinate functions on the $1$-dimensional summands. The corresponding ring $K[x_{1},x_{2},t_{1},t_{2}]^{G}$ of invariants contains the invariant $h:=(x_{1}^{2}-x_{2}^{2})t_{1}^{3}t_{2}$, separating the points $v:=([1,0]^{T},1,1)$ and $v^{\prime}:=([1,0]^{T},1,-1)$ in $K^{4}$. By Lemma 3.5, in order to prove $\beta_{\mathrm{sep}}^{K}(G)\geq 6$, it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $5$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $5$. It has positive degree in $t_{2}$, because the $(x_{1},x_{2},t_{1})$ coordinates of $v$ and $v^{\prime}$ are the same. Any $G$-invariant depending only on $t_{1}$ and $t_{2}$ is a polynomial in $t_{1}^{4}$, $t_{1}^{2}t_{2}^{2}$, $t_{2}^{4}$, and all these polynomials take the value $1$ both at $v$ and $v^{\prime}$. So $f=t_{1}^{k_{1}}t_{2}^{k_{2}}f_{1}(x_{1},x_{2})$, where $0<k_{1}+k_{2}\leq 4$, $\deg(f_{1})>0$, and $k_{2}$ is odd. Note that $t_{1}^{k_{1}}t_{2}^{k_{2}}$ is a relative $G$-invariant, hence $f_{1}(x_{1},x_{2})$ must be a relative $G$-invariant, on which $G$ acts by the inverse character. Since $c^{2}$ fixes $f_{1}$, it must fix also $t_{1}^{k_{1}}t_{2}^{k_{2}}$, so $k_{1}+k_{2}$ is even. Consequently, $(k_{1},k_{2})\in\{(1,1),(3,1),(1,3)\}$, and $f_{1}\in K[x_{1},x_{2}]$ is a relative $\mathrm{D}_{8}\cong G/\langle c^{2}\rangle$-invariant of degree at most $5-(k_{1}+k_{2})$. There is no degree $1$ relative $G$-invariant in $K[x_{1},x_{2}]$, hence $t_{1}^{k_{1}}t_{2}^{k_{2}}=t_{1}t_{2}$, and so $f$ is a relative invariant of degree at most $3$. It is easy to see that up to non-zero scalar multiples, the only relative $G$-invariants in $K[x_{1},x_{2}]_{\leq 3}$ with non-trivial weight are $x_{1}x_{2}$ and $x_{1}^{2}-x_{2}^{2}$, but none of $x_{1}x_{2}t_{1}t_{2}$ or $(x_{1}^{2}-x_{2}^{2})t_{1}t_{2}$ is a $G$-invariant. We arrived at the desired contradiction, finishing the proof. ∎
In this section
Assume that $K$ contains an element $\xi$ of multiplicative order $6$. So $\omega:=\xi^{2}$ has multiplicative order $3$, and consider the following irreducible $2$-dimensional representations of $G$:
Proposition 5.9.
We have $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2})\geq 6$.
Proof.
Note that $\ker(\psi_{1})=\langle b\rangle$ and $\ker(\psi_{2})=\langle a\rangle$. We have $G/\ker(\psi_{1})\cong\mathrm{S}_{3}$ and $G/\ker(\psi_{2})\cong\mathrm{S}_{3}$. Both these representations factor through the irreducible $2$-dimensional representation of $\mathrm{S}_{3}$. It is well known and easy to see that the corresponding algebra of invariants is generated by a degree $2$ and a degree $3$ invariant. Moreover, $K[W_{j}]/\mathcal{H}(G,W_{j})$ as an $\mathrm{S}_{3}=G/\ker(\psi_{j})$-module is isomorphic to the regular representation of $\mathrm{S}_{3}$. Thus $x_{1}^{3}-x_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $G$-module direct complement of $\mathcal{H}(G,W_{1})$ in $K[W_{1}]$ on which $\ker(\psi_{2})$ acts trivially, and $y_{1}^{3}-y_{2}^{3}$ spans the only minimal non-trivial $G$-invariant subspace in a $G$-module direct complement of $\mathcal{H}(G,W_{2})$ in $K[W_{2}]$ on which $\ker(\psi_{1})$ acts trivially. It follows that $K[W_{1}\oplus W_{2}]^{G}$ is generated by $K[W_{1}]^{G}$, $K[W_{2}]^{G}$, and $f:=(x_{1}^{3}-x_{2}^{3})(y_{1}^{3}-y_{2}^{3})$. Consider the points $v=(w_{1},w_{2})=([1,0]^{T},[1,0]^{T})$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2})=([1,0]^{T},[0,1]^{T})$. We claim that all invariants of degree at most $5$ agree on $v$ and on $v^{\prime}$, but $f(v)\neq f(v^{\prime})$. Indeed, $w_{1}=w^{\prime}_{1}$ and $c\cdot w_{2}=w^{\prime}_{2}$. Thus the invariants in $K[W_{1}]^{G}$ and in $K[W_{2}]^{G}$ agree on $v$ and on $v^{\prime}$. On the other hand, $f(v)=1$ and $f(v^{\prime})=-1$. The proof is finished. ∎
Theorem 5.10.
Proof.
By [12, Corollary 5.5] we have $\beta^{K}(G)=6$. Thus the result follows from Proposition 5.9. ∎
@@ 5.7. The group $(\mathrm{C}_{3}\times\mathrm{C}_{3})\rtimes_{-1}\mathrm{C}_{2}$
Set $v:=([1,\xi]^{T},1)\in W\oplus U_{\chi}$ and $v^{\prime}:=([1,\xi]^{T},-1)\in W\oplus U_{\chi}$. The invariant $(x_{1}^{6}-x_{2}^{6})x_{1}x_{2}t$ has different values on $v$ and $v^{\prime}$. By Proposition 5.11 we see that all $G$-invariants of degree less than $9$ agree on $v$ and $v^{\prime}$. This shows that $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\geq 9$. On the other hand, $\beta^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})=9$ by [15, Proposition 3.5], implying the reverse inequality $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\leq 9$. ∎

### v3

@@ 3.3. The binary tetrahedral group
having degrees $4,6,12,12$. In fact $h_{12}^{(1)}$ can be omitted, because we have $h_{12}^{(1)}=-4h_{6}^{2}$. Set $v:=[1,\xi^{2}]^{T}$ and $v^{\prime}:=\xi v$. Then $(x_{1}^{2}+x_{2}^{2})(v)=0=(x_{1}^{2}+x_{2}^{2})(v^{\prime})$, hence $f_{2}(v)=0=f_{2}(v^{\prime})$. It follows that $h_{6}(v)=0=h_{6}(v^{\prime})$. Moreover, $h_{8}(v)=16(\xi^{4})^{2}=16=16(\xi^{8})^{2}=h_{8}(v^{\prime})$. Thus no invariants of degree less than $12$ separate $v$ and $v^{\prime}$. However, they have different $G$-orbit, since $h_{12}^{(2)}(v)=4^{3}(\xi^{4})^{3}=-64$, whereas $h_{12}^{(2)}(v^{\prime})=4^{3}(\xi^{8})^{3}=64$. Thus we proved the inequality $\beta_{\mathrm{sep}}(G,V)\geq 12$, implying in turn $\beta_{\mathrm{sep}}^{K}(G)\geq 12$. On the other hand, by [7, Corollary 3.6] we know that $\beta^{K}(G)=12$, implying the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 12$. ∎
@@
Set $v:=([1,\xi]^{T},1)\in W\oplus U_{\chi}$ and $v^{\prime}:=([1,\xi]^{T},-1)\in W\oplus U_{\chi}$. The invariant $(x_{1}^{6}-x_{2}^{6})x_{1}x_{2}t$ has different values on $v$ and $v^{\prime}$. By Proposition 3.6 we see that all $G$-invariants of degree less than 99 agree on $v$ and $v^{\prime}$. This shows that $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\geq 9$. On the other hand, $\beta^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})=9$ by [11, Proposition 3.5], implying the reverse inequality $\beta_{\mathrm{sep}}^{K}((\mathrm{C}_{6}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{2})\leq 9$. ∎
3.5. The symmetric group $\mathrm{S}_{4}$ of degree $4$.
Theorem 3.8.
Proof.
We have $\beta^{K}(\mathrm{S}_{4})=9$ by [10, Example 5.3], implying the inequality $\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{4})\leq 9$.
In order to show the reverse inequality consider $V=K^{4}$ endowed with the tensor product of the sign representation of $\mathrm{S}_{4}$ and the standard action of $\mathrm{S}_{4}$ on $K^{4}$ via permuting the coordinates. Thus $\pi\in\mathrm{S}_{4}$ maps the variable $x_{j}$ to $\mathrm{sign}(\pi)x_{\pi(j)}$. Denote by $\sigma_{j}$ ($j=1,2,3,4)$ the elementary symmetric polynomial of degree $j$, and set $\Delta:=\prod_{1\leq i<j\leq 4}(x_{i}-x_{j})$. The algebra $K[V]^{\mathrm{S}_{4}}$ is minimally generated by $\sigma_{1}^{2}$, $\sigma_{2}$, $\sigma_{1}\sigma_{3}$, $\sigma_{4}$, $\sigma_{3}^{2}$, $\sigma_{1}\Delta$, $\sigma_{3}\Delta$ (see [10, Example 5.3]).
We shall show that there exists a $v\in V$ with
$(x^{2}+x+1)(x^{2}-x+2)=x^{4}+2x^{2}+x+2\in K[x],$
3.6. The Heisenberg group $\mathrm{H}_{27}$
In this subsection
Write $V$ for the vector space $K^{3}$ endowed with the representation $\psi$.
Proposition 3.9.
Proof.
Denote by $H$ the subgroup of $G$ generated by $a$ and $c$. The algebra $K[V]^{H}$ is generated by the monomials $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$, $x_{1}x_{2}x_{3}$. It follows that each homogeneous $G$-invariant has degree divisible by $3$. The element $b$ fixes $x_{1}x_{2}x_{3}$, and permutes cyclically $x_{1}^{3}$, $x_{2}^{3}$, $x_{3}^{3}$. So $K[V]^{G}$ is generated by $x_{1}x_{2}x_{3}$ and the generators of $K[x_{1}^{3},x_{2}^{3},x_{3}^{3}]^{\langle b\rangle}$. One easily deduces that the $\langle b\rangle$-invariants in $K[V]^{H}$ (i.e. the $G$-invariants in $K[V]$) of degree at most $6$ are contained in the subalgebra generated by $f_{1}:=x_{1}x_{2}x_{3}$, $f_{2}:=x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$, $f_{3}:=x_{1}^{3}x_{2}^{3}+x_{1}^{3}x_{3}^{3}+x_{2}^{3}x_{3}^{3}$. If $|K|\neq 4$ then there exists a non-zero $\lambda\in K$ with $\lambda^{3}\neq 1$. Now $f_{1}$, $f_{2}$, and $f_{3}$ are all symmetric in the variables $x_{1},x_{2},x_{3}$, therefore they agree on $v:=[1,\lambda,0]^{T}$ and $v^{\prime}:=[\lambda,1,0]^{T}$. On the other hand, $v$ and $v^{\prime}$ have different $G$-orbits. Indeed, the algebra $K[V]^{G}$ contains also $f_{4}:=x_{1}^{3}x_{2}^{6}+x_{1}^{6}x_{3}^{3}+x_{2}^{3}x_{3}^{6}$. We have $f_{4}(v)=\lambda^{6}$, whereas $f_{4}(v^{\prime})=\lambda^{3}$, so $f_{4}(v)\neq f_{4}(v^{\prime})$. Thus $G\cdot v\neq G\cdot v^{\prime}$, and $v$, $v^{\prime}$ can not be separated by invariants of degree <9<9.
Theorem 3.10.
Proof.
By [6, Corollary 18] we have $\beta^{K}(G)=9$, therefore the result follows by Proposition 3.9. ∎